OpenStudy (anonymous):

1/2+(2/3 / 3/4)-(4/5 x 5/6)=?

OpenStudy (anonymous):

1/2+1/2-2/3===1/3

OpenStudy (anonymous):

helped

OpenStudy (imer):

\[\frac{ 1 }{ 2 }+(\frac{ \frac{ 2 }{ 3 } }{ \frac{ 3 }{ 4 } })-(\frac{ 4 }{ 5 }*\frac{ 5 }{ 6 }) \rightarrow \frac{ 1 }{ 2 }+(\frac{ 2(4) }{ 3(3) }) -(\frac{ 4 }{ 6 }) \rightarrow \frac{ 13 }{ 18}\]

OpenStudy (imer):

@kritikachandak, you answer is wrong.

OpenStudy (anonymous):

1/2+8/9-2/3=(9+16+12)/18=37/18=\[2\left(\begin{matrix}1 \\ 18\end{matrix}\right)\]

OpenStudy (anonymous):

sorry (25-12)/18=13/18

OpenStudy (anonymous):

sorry (25-12)/18=13/18